### v1

# Fast and Accurate Charge Transfer Excitations via Nested Aufbau Suppressed Coupled Cluster

Abstract
Modeling charge transfer well can require treating post-excitation orbital relaxations and handling medium to large molecules in realistic environments. By combining a state-specific correlation treatment with such orbital relaxations, Aufbau suppressed coupled cluster has proven accurate for charge transfer, but, like many coupled cluster methods, it struggles with large system sizes. We derive a low-cost Aufbau suppressed second order perturbation theory and show that, by nesting a small coupled cluster treatment inside of it, computational cost and scaling are reduced while accuracy is maintained. Formal asymptotic costs are dropped from iterative $N^{6}$ to non-iterative $N^{5}$ plus iterative $N^{3}$, and we test an initial implementation that can handle about 100 atoms and 800 orbitals on a single computational node. Charge transfer excitation energy errors are below 0.1 eV on average, a 0.25 eV improvement over $N^{6}$-cost equation of motion coupled cluster with singles and doubles.
preprint: AIP/123-QED
Charge transfer (CT) excitations are at the crux of photochemical efficacy across a wide range of current scientific interests, including photovoltaics,[1] photosynthetic complexes,[2] and photocatalysis.[3] However, traditional linear response approaches such as time dependent density functional theory (TD-DFT) [4, 5, 6] and equation of motion coupled cluster (EOM-CCSD) [7, 8, 9] produce significant errors for these excitations [10, 11, 12, 13, 14, 15] due in part to their limited ability to capture nonlinear orbital relaxation effects. [16, 17] The challenge of modeling CT is further exacerbated by system size: these excitations often occur in medium to large donor-bridge-acceptor molecules [18] and can be strongly influenced by solvent and other environmental effects. [19] To model charge transfer well, a method must capture orbital relaxations and their effects on electron correlation while maintaining a low computational cost. Our recently introduced Aufbau suppressed coupled cluster (ASCC) approach [20, 21] captures both orbital relaxation and its correlation effects, but, with a cost like the ground state singles and doubles theory (CCSD), its applicability to CT was initially somewhat limited. In this study, we develop a second order perturbative counterpart to ASCC and, by nesting a small coupled cluster evaluation inside of it, reduce the cost and scaling of the approach while maintaining its accuracy for valence, Rydberg, and CT excitations.
ASCC and its partially linearized cousin (PLASCC) have a two-fold advantage in handling post-excitation orbital relaxations that makes them particularly well positioned to model CT. [20, 21] First, they can be constructed on top of a reference wave function that already captures orbital relaxation effects at the mean field level. Second, their full $\hat{T}_{1}$ singles operators allow them to fine tune orbital relaxations in the presence of correlation effects.[22] The ability to deliver a CCSD-like weak correlation treatment alongside and informed by these orbital relaxations has allowed PLASCC to deliver $\sim$0.1 eV excitation energy errors across a wide variety of excited states, including Rydberg and CT states. [21] While this accuracy is especially noteworthy for CT, where EOM-CCSD’s errors are $\sim$0.35 eV, the $N^{6}$ cost scaling with system size $N$ has so far limited the applicability of ASCC and PLASCC to relatively small systems.
To lower cost while maintaining accuracy, we propose to exploit ASCC theory’s strong connections to perturbation theory (PT) to avoid wasting effort perfecting the treatment of correlation effects that are simply going to cancel out when evaluating the excitation energy difference. Just as ground state CCSD is motivated by PT and intimately linked to second order Møller–Plesset (MP2) theory, [23, 24, 25, 26, 27] ASCC and especially PLASCC are motivated by PT arguments [21] and, as we will show, share an analogous link to an excited state second order Aufbau suppressed PT. Exploiting this link, we will construct a nested ASCC approach in which the low cost PT determines which orbitals’ correlation effects are most strongly influenced by the excitation, and which orbitals’ correlations are sufficiently insensitive to the excitation that leaving their treatment at the PT level should be safe. A subsequent ASCC or PLASCC treatment of only the strongly affected orbitals will then be used to refine the correlation effects that most strongly influence the excitation energy.
In order to arrive at a second order excited-state-specific PT, we will begin with the ASCC equations and systematically remove terms contributing at higher perturbative orders, similarly to how MP2 can be derived from ground state CCSD.[23] The ASCC ansatz is exponentially parameterized using both an excitation operator $\hat{T}$ and a one-body deexcitation operator $\hat{S}^{\dagger}$ that act on a closed-shell Aufbau determinant, |$\ket{\phi_{0}}$.[20]
|$\ket{\Psi_{\text{ASCC}}}=e^{-\hat{S}^{\dagger}}e^{\hat{T}}\ket{\phi_{0}}$ (1)
With this choice of ansatz, remarkably simple initializations of $\hat{T}$ and $\hat{S^{\dagger}}$ allow us to transform |$\ket{\phi_{0}}$ into a qualitatively correct open-shell excited state, at which point $\hat{T}$ can be further optimized to capture weak correlation effects.[20] We move towards working equations in the same manner as traditional CC, which leads to similarity transformations of the Hamiltonian.
$\bar{H}$ $=e^{-\hat{T}}e^{\hat{S}^{\dagger}}\hat{H}e^{-\hat{S}^{\dagger}}e^{\hat{T}}$ (2)
Since $\hat{S}^{\dagger}$ is a one-body operator, the inner transformation amounts to a two-electron integral transform. Working with the outer transformation leads to tensor contractions similar to those in CCSD, but which are evaluated using the modified integrals resulting from the inner transform. As in ground state CC, the ASCC energy is evaluated as the expectation value with respect to |$\ket{\phi_{0}}$, while projections with excited determinants $\bra{\phi_{\mu}}$ | yield the amplitude equations that we solve to optimize $\hat{T}$.
$E$ $=\bra{\phi_{0}}\bar{H}\ket{\phi_{0}}$ (3)
00 0 $=\bra{\phi_{\mu}}\bar{H}\ket{\phi_{0}}$ (4)
Considering these equations now from a perturbative perspective, the initialization that both $\hat{S}^{\dagger}$ and $\hat{T}$ require to arrive at a qualitatively correct starting point implies that they both contain nontrivial zeroth order pieces. [21] Note, however, that $\hat{S}^{\dagger}$ remains fixed after initialization, with only $\hat{T}$ varying and thus containing higher order contributions.
$\hat{S}^{\dagger}$ $=\hat{S}^{\dagger(0)}$ (5)
$\hat{T}$ $=\hat{T}^{(0)}+\hat{T}^{(1)}+\hat{T}^{(2)}+\cdots$ (6)
After perturbatively partitioning the Hamiltonian (see Supporting Information), its double similarity transform can also be partitioned by perturbative order.
$\hat{H}$ $=\hat{H}^{(0)}+\hat{H}^{(1)}$ (7)
$\bar{H}$ $=\bar{H}^{(0)}+\bar{H}^{(1)}+\bar{H}^{(2)}+\cdots$ (8)
Substituting this expansion of $\bar{H}$ into the energy and residual equations (Eqns. 3 and 4) allows us to evaluate cluster amplitude and energy contributions order by order. By adding together all energetic contributions up to second order, we are thus able to determine the cluster amplitudes necessary to achieve a second order excited-state-specific energy, thus arriving at the second order PT we will need for our nested approach.
$E_{\text{PT2}}$ $=E^{(0)}+E^{(1)}+E^{(2)}$ (9)
$E^{(n)}$ $=\bra{\phi_{0}}\bar{H}^{(n)}\ket{\phi_{0}}$ (10)
00 0 $=\bra{\phi_{\mu}}\bar{H}^{(n)}\ket{\phi_{0}}\rightarrow\hat{T}_{\mu}^{(n)}$ (11)
While the working equations for this PT are a bit more complicated than for MP2 due to the presence of $\hat{T}^{(0)}$, they do closely mirror MP2’s computational cost. Once the two-electron integrals are available in the molecular orbital (MO) basis, every term in the working equations scales as $O(N^{4})$ or less, as is the case for MP2. Even better, the amplitude equations in Eqn. 11 are linear in the highest order terms — which at each order are the ones being solved for — and are small-block block-diagonal within a semicanonical basis, allowing them to be solved in a non-iterative manner (see Supporting Information). Therefore, the headline scaling of this PT, in exact asymptotic equivalence to MP2, comes from the non-iterative $O(N^{5})$ cost of transforming the two-electron integrals into the MO basis. Compared to the iterative $O(N^{6})$ scaling of ASCCSD, this is a substantial improvement in computational efficiency and notably avoids some of the most memory intensive terms found in CCSD and ASCCSD. Even when compared to relatively affordable wave function correlation methods for excited states, such as the iterative $O(N^{5})$ scaling CC2[28] and ADC(2)[29, 30, 31] methods, the non-iterative $O(N^{5})$ scaling of a routine integral transformation represents a significant cost reduction.
Figure 1: Excitation energy error distributions for various methods on 130 valence and Rydberg single-CSF singlet states from the QUEST benchmark in an aug-cc-pVDZ basis. For the PT alone, 17 outlier states with a $<-1.0$ eV error are not shown. Note the change of scale in the y-axis between the top and bottom rows. The mean signed error (MSE), mean unsigned error (MUE), and standard deviations (SD) are shown in insets. Reference values are of at least of EOM-CCSDT quality.
Figure 2: Mean unsigned errors (MUEs) for various methods on 7 CT excitations coming mostly from the benchmark of Szalay and coworkers in a cc-pVDZ basis. MUEs extending beyond the plot are printed at the top of each bar. Reference values are EOM-CCSDT except for one LR-CC3 value.
Before we go about nesting a small ASCC treatment within this PT, it is worthwhile to evaluate its accuracy as a standalone method. To do so, we test the PT on a collection of 130 singly excited, single CSF, singlet valence and Rydberg states from the QUEST molecular benchmark[32, 33] as well as a set of 7 CT states mainly originating from the benchmark of Szalay and coworkers.[14] The results are summarized in Figures 1 and 2, with additional details available in the Supporting Information. While computationally efficient, the PT is unsurprisingly less accurate than its parent CC theory. Though it performs well in many cases, it produces a substantial number of severe errors, some larger than an eV, particularly in the aromatic molecules where PLASCC makes the largest improvements over ASCC. Indeed, this pattern helps corroborate the original motivation for PLASCC, which was that in these systems the perturbation is large enough that higher order and poorly balanced terms in ASCC become important. [21] At the very least, these PT results suggest that the perturbation in ASCC theory is indeed larger in many of these ring states.
Of particular concern for the present study is that the PT fares worse in CT states (errors of $\sim$0.6 eV) than it does in valence and Rydberg states (errors of $\sim$0.4 eV). This worsening contrasts sharply with PLASCC, which produces errors of $\sim$0.1 eV in each category. One culprit is likely the truncation of singles contributions to low orders in PT theory, which interferes with the parent theory’s ability to provide CC-singles-style corrections to the initial mean field orbital relaxations. [20, 21] Another factor is that more electron pairs experience large, excitation-induced changes to their electron correlation in CT states, because CT excitations make major changes to how close various electrons are to each other. It is to be expected that second order PT would be less effective at capturing these changes than a CCSD-style approach. These challenges notwithstanding, the PT does offer significantly higher accuracy for CT than hybrid TD-DFT methods, even when long-range-corrected functionals like $\omega$B97X-D [34] and CAM-B3LYP [35] are used.
To return accuracy to the CC level while keeping cost in check, we now turn to nesting a small CC treatment of the most important correlations within a larger PT treatment of the rest. We begin by noting that, even in CT excitations that move electrons around, the number of individual orbitals that are directly involved in or strongly affected by the excitation is not likely to grow with system size, and that these orbitals tend to be spatially localized (e.g. on the donor and acceptor). This motif suggests that, at least in a large molecule, the vast majority of the correlation contributions will come from orbitals only weakly influenced by the excitation. We might therefore expect such contributions to be about the same in the ground and excited state and to mostly cancel out in the excitation energy difference. If that is indeed the case, it is likely unnecessary to treat all correlations at the CC level. Instead, we propose that PT should be sufficient for most of them while a more targeted CC treatment handles the few that are strongly affected by the excitation.
We therefore begin by performing full MP2 and Aufbau suppressed PT calculations on the ground and excited states, respectively. While these use (semi)canonical bases to ensure a non-iterative $O(N^{5})$ cost, we then proceed to localize and match orbitals between the ground and excited state (see Supporting Information) in preparation for the correlation analysis. In this matched local basis, we construct ground and excited state measures for the correlation contributions related to each orbital. Specifically, for a given orbital, we add together all of the PT’s amplitude contributions to Eq. 9 in which the amplitude involved has at least one index on that orbital. We then compare this correlation measure between the ground and the excited state, and, if they differ by more than a small set threshold, that orbital gets flagged as needing to be part of the nested CC treatment. One can imagine other screening approaches, and it’s hard to know what the optimal approach would be, but we will emphasize some strong formal properties of this approach: it preserves size consistency, extensivity, and intensivity and, for a given localized orbital, is invariant to rotations among the other localized occupied or virtual orbitals.
Having identified the orbitals whose correlation is strongly affected by the excitation, we are ready to perform a small ASCC refinement nestled within the larger PT treatment. To do so, we freeze at the PT level all amplitudes with one or more indices on weakly affected orbitals and proceed to solve the ASCC or PLASCC equations for the remaining amplitudes, using their PT values as an initial guess. We therefore only need to work with the amplitude residuals (Eq. 4) for which all residual indices are on strongly affected orbitals. Once we have solved these residual equations for their corresponding amplitudes, we take these refined amplitudes along with those frozen at the PT level and perform one final whole-system evaluation of the CC energy in the ground and excited states.
As there are expected to be only $O(1)$ orbitals whose correlation is strongly affected by the excitation, this nested approach leads to a dramatic cost reduction compared to the iterative $O(N^{6})$ cost of a full ASCC evaluation. We can break the cost of the nested CC into two parts: the non-iterative part and the iterative part. The final energy evaluation, as in standard CCSD, carries a non-iterative $O(N^{4})$ cost, and indeed all non-iterative portions of the nested CC are $O(N^{4})$ or better. To see why, note that amplitude indices ranging over $O(N)$ values must contract with the Hamiltonian if their amplitude is to contribute to one of the $O(1)$ residuals we are evaluating. As the Hamiltonian remains two-body even after the $\hat{S}^{\dagger}$ transformation, it has only four indices to work with, and so any such term can have no worse than an $O(N^{4})$ cost scaling.
Similar arguments reveal that the iterative portion of the nested CC need not have a cost worse than $O(N^{3})$. We note that, due to the connectedness of the CC equations, any amplitude appearing in the residual equations must contract with the Hamiltonian. If all four of the Hamiltonian’s indices participate in $O(N)$-range contractions, then we know that none of the amplitudes are among those updated in the nested treatment, and so the contribution in question can be evaluated once and stored, becoming one of the non-iterative terms. Thus, there cannot be more than three indices with $O(N)$ ranges in each term that actually changes during the iterative solution of the amplitude equations, and so we conclude that the iterative cost is not worse than $O(N^{3})$. While a careful use of intermediates may be able to drop this iterative scaling further, in the present study we have focused on ensuring that our implementation achieves this iterative $O(N^{3})$ bound. Thus, all told, our nested CC implementation involves non-iterative $O(N^{4})$ steps and iterative $O(N^{3})$ steps, while the transformation of the PT amplitudes into the local basis for the correlation analysis involves a non-iterative $O(N^{5})$ step, much like a standard two-electron integral transform would.
We now turn to the question of whether the nested approach maintains the accuracy of its parent CC theory. Figure 1 reveals that, for valence and Rydberg states, the nested approach is far more accurate than the PT alone and, at least for PLASCC, has an accuracy commensurate with the full CC treatment. Interestingly, nested ASCC is more accurate than full ASCC, which is likely driven by the same effect that makes PLASCC more accurate than ASCC. We observed previously [21] that certain nonlinear terms in ASCC are less well balanced than in the ground state due to the perturbative incompleteness of their energy contributions. PLASCC counteracts this effect through a partial linearization that explicitly deletes the offending terms, whereas nested ASCC appears to at least partially counteract it by freezing many of the amplitudes at a specific order in PT theory, where such imbalances do not exist because all terms contributing at a given order are included. As we saw when testing PLASCC, [21] this effect can be observed clearly in nested ASCC’s tendency to reduce symmetry breaking relative to full ASCC. In addition to good accuracy, these tests also reveal that, even in the modestly sized molecules in this benchmark, the cost reduction from nesting is significant. While the number of occupied orbitals whose correlation contributions are frozen at the PT level in these small molecules is roughly what one would expect from the frozen core approximation, the number of virtual orbitals with PT-frozen correlation is quite substantial, typically 40% or more of the virtual set.
Turning now to the CT results in Figure 2, we see a similarly encouraging story unfold. Nested ASCC and PLASCC are much more accurate than the PT theory on its own. PLASCC and nested PLASCC are again more accurate than ASCC, although in this case nested ASCC is not an improvement over full ASCC. Comparing to other excited state theories, we see that nested PLASCC maintains PLASCC’s 0.25 eV improvement over EOM-CCSD, which is now especially significant in light of the non-iterative $O(N^{5})$ vs iterative $O(N^{6})$ computational costs. Nested PLASCC far outstrips the TD-DFT accuracy of the density functionals we test here, and is also a notable improvement over the much wider range of TD-DFT methods tested by Mester and Kállay. [13] Seeing that PLASCC’s accuracy is maintained under nesting in valence, Rydberg, and even CT states, we now turn our attention to evaluating computational efficiency in larger systems where tidy theoretical energy benchmarks are not available.
Figure 3: A gas phase, simple valence excitation (A), a simple valence excitation in a solvated environment (B), and a gas phase, donor-bridge-acceptor CT excitation (C) computed with the nested ASCC approach. Pink surfaces are 95% isosurfaces for the electronic density of the occupied orbitals receiving a CC correlation treatment, as automatically determined by the PT’s correlation analysis.
We begin with the simple gas phase valence excitation in Figure 3A to demonstrate a few key points. First, this excitation is an example where the nested CC refinement makes a significant, $\sim$0.15 eV difference compared to the PT alone. This difference is likely driven by the importance of post-excitation orbital relaxations on the relatively polarizable sulfur atom following the delocalization of the lone pair electron into the $\pi^{*}$ system. Second, despite this delocalization, the excitation remains localized compared to the extent of the molecule, which should leave most orbitals’ correlation contributions mostly unaffected. Finally, its computational cost grows with system size as we would expect it to, as seen in Figure 4.
Figure 4: Wall time measured for Aufbau Suppressed PT (top), the noniterative step in nested ASCCSD (middle), and the iterative step in nested ASCCSD (bottom) on one Intel Xeon Gold 5218 processor containing 32 cores for the $n$ to $\pi^{*}$ excitation in thiopropynal with a varying length alkane chain. Dashed lines show best fit curves of the form $y\propto x^{a}$ fit to the last 10 data points.
For an explicit solvation example, we examine in Figure 3B the $n$ to $\pi^{*}$ transition in solvated acetone. The explicit waters play a non-innocent role, as the hydrogen bonding that initially stabilizes the ground state is disrupted when one of the hydrogen-bond-accepting lone pair electrons is transferred into the $\pi^{*}$ orbital by the excitation. Crucially, despite the strong solute-solvent interactions and the three-dimensional nature of this example, the orbitals flagged for CC refinement remain few in number and are tightly localized in space. We find it especially noteworthy that, despite no explicit input to this effect from the user, the PT’s correlation analysis selected CC orbitals in a physically intuitive manner: the water orbitals it picks out are the ones that had their strong hydrogen bonding interactions perturbed by the excitation.
Finally, we examine the donor-bridge-acceptor CT example of Figure 3C, which results from adding an alkane bridge to one of our earlier CT tests. In contrast to the last two examples in which the excitation was localized in only one region of space, here the PT’s correlation analysis flags two separate regions for CC treatment: one on the donor and one on the acceptor, again following what physical intuition would suggest. While this donor/acceptor effect leads to a larger number of orbitals in the CC calculation, this number stops changing with system size once the bridge reaches just three carbons in length. Thus, regardless of whether the excitation itself is a local valence excitation or a non-local CT, we see that the nesting approach successfully demarcates a physically intuitive set of orbitals for careful CC treatment while dramatically reducing the computational cost compared to a full-system CC evaluation.
In conclusion, we have developed an approach for replicating the high accuracy of excited-state-specific Aufbau suppressed coupled cluster at low cost by nesting a small coupled cluster evaluation within a larger perturbative treatment. This new perturbation theory displays strong parallels to MP2, including the headline $N^{5}$ cost scaling from its bottleneck integral transformation and its strong connections to coupled cluster theory. We find it especially interesting that a simple analysis of this perturbation theory’s correlation contributions is able to automatically demarcate physically intuitive orbital subsets for which a coupled-cluster-level treatment is most essential when predicting excitation energies. While the perturbation theory is not particularly accurate on its own, the overall nesting approach rivals the parent PLASCC method in accuracy in valence, Rydberg, and CT states. This accuracy is especially promising given the approach’s headline cost reduction from iterative $N^{6}$ to non-iterative $N^{5}$ and the dramatically larger system sizes that can be reached by an initial implementation.
While this nesting approach is a significant step forward, it is only one step in a larger journey towards high-accuracy, low-cost models of charge transfer excitations. Given the strong parallels between the new perturbation theory and MP2, as well as those between ASCC and CCSD, it seems likely that local correlation techniques will further reduce both the scaling and cost of this approach. Accuracy can likely be improved as well, as the same perturbative framework used here should also be effective for deriving higher-order non-iterative perturbative corrections to the ASCC energy, as has been so richly explored in the ground state. These continuing advances promise to bring ever larger molecules and more complicated chemical environments within reach of high-accuracy excited state modeling.
I Computational Methods
Our EOM-CCSD calculations and Foster-Boys orbital localizations[36] were performed with PySCF,[37, 38, 39] our DFT and EOM-CCSDT calculations were performed with Q-Chem,[40] and our LR-CC3 calculations were performed with PSI4.[41] Though the literature and reference calculations were performed utilizing the frozen core approximation, our EOM-CCSD, PT, and ASCC calculations were performed without this approximation. Nevertheless, this is expected to make only a small difference to the excitation energies ($\sim$0.02 eV)[32, 33] and thus the high accuracy frozen-core values still make excellent references. The geometries for the large molecule valence and charge transfer excitations were determined via MP2/cc-pVDZ geometry optimizations in Q-Chem, while the solvated acetone geometry was determined using the solvator method in ORCA.[42, 43] Explicit geometries can be found in the Supporting Information.
For the excited state PT, ASCC, and PLASCC calculations, the excited state reference was determined via the excited state mean field (ESMF) method,[44, 45, 46] which can be summarized as a CIS wavefunction with full orbital relaxations. Any CSF in the ESMF reference with a singular value greater than 0.2 was considered part of the excited state reference. Block-sparse [47] Cholesky decompositions [48] with a threshold of $10^{-7}$ Hartree were used to evaluate MO basis two electron integrals, with Libint[49] supplying the underlying atomic orbital integrals. For the ASCC based approaches, convergence was considered achieved when the maximum residual was no greater than $10^{-5}$ and the energy changed by less than $10^{-7}$ Hartree in an iteration. For all other methods, default convergence criteria were used. On rare occasions, convergence difficulties were encountered for ASCC based approaches. Details regarding these states may be found in the Supporting Information. For the nested ASCC calculations, three orbital energy thresholds for establishing the small orbital set for the CC part were tested. The accuracy data in the main text are for a 0.005 eV threshold, with additional accuracy data for 0.01 eV and 0.0025 eV thresholds available in the SI. The scaling analysis was performed using the 0.01 eV threshold.
II Acknowledgements
This work was supported by the National Science Foundation, Award Number 2320936. Calculations were performed using the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory. H.T. acknowledges that this material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
S1 Perturbatively Partitioning the Hamiltonian
In order to utilize perturbation theory, a choice must be made as to how to partition the Hamiltonian. In the ground state MP theory with a general single determinant (not necessarily Hartree Fock), the zeroth order Hamiltonian is chosen to be the occupied-occupied and virtual-virtual blocks of the one-electron Fock operator while the first order Hamiltonian is everything else.
$\hat{H}$ $=H^{(0)}+H^{(1)}$ (S1)
$\hat{H}^{(0)}$ $=\sum_{ij}f_{ij}\hat{i}^{\dagger}\hat{j}+\sum_{ab}f_{ab}\hat{a}^{\dagger}\hat{b}$ (S2)
$\hat{H}^{(1)}$ $=\hat{H}-\hat{H}^{(0)}$ (S3)
In the Aufbau Suppressed PT used in this study, an additional further subdivision must be made in order to delineate the importance of the primary hole and particle orbitals being excited from and into during a particular excitation. Thus, we will now take the indices $ij\cdots$ and $ab\cdots$ to represent nonprimary occupied and virtual orbitals, respectively. $h$ and $p$ will refer to hole and particle orbitals, respectively.
$\hat{H}$ $=H^{(0)}+H^{(1)}$ (S4)
$\hat{H}^{(0)}$ $=\sum_{h_{1}h_{2}}f_{h_{1}h_{2}}\hat{h}_{1}^{\dagger}\hat{h}_{2}+\sum_{ij}f_{ij}\hat{i}^{\dagger}\hat{j}$
$\quad+\sum_{p_{1}p_{2}}f_{p_{1}p_{2}}\hat{p}_{1}^{\dagger}\hat{p}_{2}+\sum_{ab}f_{ab}\hat{a}^{\dagger}\hat{b}$ (S5)
$\hat{H}^{(1)}$ $=\hat{H}-\hat{H}^{(0)}$ (S6)
Though slightly more complicated, this form of $\hat{H}^{(0)}$ maintains a block diagonal structure and ensures that the only zeroth order cluster amplitudes are those mentioned in the main text: the subset of amplitudes initialized to nonzero values during the Aufbau-suppression-style construction of the qualitatively correct starting point. When this starting point is a single CSF, as was true for all states in this study, the hole and particle sums in Eq. S5 each contain just one term, because in these cases there are just one hole orbital and one particle orbital. In order to maintain equivalence between the perturbation theories of the ground and excited states, this partitioning scheme was also utilized for the ground state, which causes the ground state energy to differ slightly from the usual MP2 energy. However, this choice does slightly improve the PT’s accuracy, as can be seen in Figure S1.
Figure S1: Error distributions for the Aufbau Suppressed PT’s excitation energy predictions on the QUEST set when using either the altered zeroth order Hamiltonian (left) or the traditional MP2 choice (right) in the ground state calculation. Mean signed errors (MSEs), mean unsigned errors (MUEs), and standard deviations (SD) are shown for both methods.
S2 Solution to the PT Equations
In the canonical orbital basis, the diagonality of the Fock operator allows the MP2 linear equation for the amplitudes to be solved in a non-iterative manner, as it guarantees that each linear equation is fully decoupled from the others. In our PT, while many of the linear equations are also decoupled in the same way when expressed in a semicanonical basis, the presence of the off-diagonal, deexciting zeroth order Hamiltonian elements in conjunction with zeroth order cluster amplitudes causes some of the linear equations to be coupled. As an example of these coupled equations, the highest order perturbative amplitude terms are shown for the all-primary singles and doubles residual equations in the semicanonical basis for the single-CSF case.
$\bar{H}^{(0)}$ $=e^{\hat{S}^{\dagger}}\hat{H}^{(0)}e^{-\hat{S}^{\dagger}}$
$=\sum_{h_{1}h_{2}}\bar{f}_{h_{1}h_{2}}\hat{h}_{1}^{\dagger}\hat{h}_{2}+\sum_{ij}\bar{f}_{ij}\hat{i}^{\dagger}\hat{j}$
$\quad+\sum_{p_{1}p_{2}}\bar{f}_{p_{1}p_{2}}\hat{p}_{1}^{\dagger}\hat{p}_{2}+\sum_{ab}\bar{f}_{ab}\hat{a}^{\dagger}\hat{b}$
$\quad+\sum_{hp}\bar{f}_{hp}\hat{h}^{\dagger}\hat{p}$ (S7)
${R_{h}^{p}}^{(n)}$ $=\left(\bar{f}^{(0)}_{pp}-\bar{f}^{(0)}_{hh}-2\bar{f}_{hp}^{(0)}{t_{h}^{p}}^{(0)}\right){t_{h}^{p}}^{(n)}+\bar{f}_{\bar{h}\bar{p}}^{(0)}{t_{h\bar{h}}^{p\bar{p}}}^{(n)}+\cdots$ (S8)
${R_{h\bar{h}}^{p\bar{p}}}^{(n)}$ $=\left(-2\bar{f}_{hp}^{(0)}{t_{h}^{p}}^{(n)}-2\bar{f}^{(0)}_{\bar{h}\bar{p}}{t_{\bar{h}}^{\bar{p}}}^{(n)}\right){t_{h\bar{h}}^{p\bar{p}}}^{(0)}+\left(\bar{f}_{pp}^{(0)}+\bar{f}_{\bar{p}\bar{p}}^{(0)}-\bar{f}_{hh}^{(0)}-\bar{f}_{\bar{h}\bar{h}}^{(0)}-2\bar{f}_{hp}^{(0)}{t_{h}^{p}}^{(0)}-2\bar{f}_{\bar{h}\bar{p}}^{(0)}{t_{\bar{h}}^{\bar{p}}}^{(0)}\right){t_{h\bar{h}}^{p\bar{p}}}^{(n)}+\cdots$ (S9)
If the number of equations in a coupled block scaled with system size, a non-iterative solution would become prohibitively expensive, and iterative methods would be required. However, because the off-diagonal zeroth order Hamiltonian element deexcites only in the small primary orbital space, the blocks of coupled equations are small and independent of system size. In the example shown above, because $\bar{H}^{(0)}$ contains exactly one unique off-diagonal element, only two linear equations are coupled. In the single-CSF case, the largest blocks contain just 6 coupled linear equations. Thus, solving the PT’s amplitude equations amounts to explicitly inverting $O(N^{4})$ small matrices, none of which have dimensions that grow with system size. Although a bit more complicated than the fully diagonal MP2 equations, the cost of solving our PT’s equations is also non-iterative $O(N^{4})$. Thus, as for MP2, the bottleneck in large systems is actually the $O(N^{5})$ cost of transforming the two-electron integrals into the molecular orbital basis.
S3 Orbital Transformations for Nested Coupled Cluster
Upon completion of the perturbation theory calculations, the orbitals are in their respective semicanonical bases for both the ground and excited states, which may differ due to post-excitation orbital relaxations. However, in order to exploit locality when determining which correlations are and are not strongly affected by the excitation, we prefer to transform from the semicanonical basis into a local orbital basis. To do so, the excited state’s non-hole occupied and non-particle virtual orbitals are separately localized using the Foster-Boys method. Note that we leave the hole and particle orbitals unchanged in order to maintain the compactness of the determinant representation of the excited state. At this point, we construct matching ground state orbitals by projecting each of the excited state’s (hole + localized) occupied orbitals into the span of the ground occupieds, after which a Löwdin orthonormalization provides an orthonormal set of ground state occupied orbitals whose shapes are closely matched to the excited state occupieds. We repeat this process for the (particle + localized) excited state virtual orbitals to produce a similarly matched ground state virtual basis.
Now that we have localized ground and excited state orbitals whose shapes and therefore physical interpretations are similar, we perform the correlation analysis described in the main text that separates the molecular orbitals into two sets: those that will have their correlation frozen at the PT level and those that will have it refined via CC. Note that, to perform this analysis, we need to first rotate the PT amplitudes into the local basis, which incurs a non-iterative $O(N^{5})$ cost similar to that of a standard two electron integral transformation. Finally, to improve the approximate Jacobian for the iterative solution of the nested CC treatment, the set of non-hole and non-particle orbitals flagged for CC treatment are returned to a semicanonical basis, again leaving the hole and particle untouched to maintain the compact form of the excited state. The PT2 amplitudes are rotated into this new basis as well, this time at just an $O(N^{4})$ cost thanks to savings related to the $O(1)$ range of indices for orbitals receiving the CC refinement.
S4 QUEST and Charge Transfer Benchmark
The full table of results for the QUEST and charge transfer benchmarks may be found in a .xlsx file available upon request. On rare occasion, an ASCC-based method would experience convergence difficulties, particularly when the ESMF wave function had significant deficiencies as compared to EOM-CCSD. If neither of the two solutions would converge for any of the ASCC-based methods, they were entirely removed for all methods. States where only one of the solutions failed to converge or only a subset of methods failed to converge are highlighted in red. For the states where only one of two solutions would converge, the final reported energy is not the arithmetic mean of the two solutions but just the energy of the only converged solution. Finally, a handful of states would stall near but not quite at our tight convergence criterion. These states, highlighted in blue, are still included due to their energy changing below the level of energetic precision reported in this work.
Three different energetic thresholds for the nesting procedure were tested in this study: 0.0025 eV, 0.005 eV, and 0.01 eV. Ultimately, we decided to report the results of the 0.005 eV threshold in the main text due to its balance of computational efficiency and accuracy. The accuracy for each of the three tested thresholds for the QUEST benchmark are summarized in Figure S2.
Figure S2: Excitation energy error distributions for ASCCSD (left) and PLASCCSD (right) with differing nesting orbital energy thresholds in eV on 130 valence and Rydberg single-CSF singlet states from the QUEST benchmark in an aug-cc-pVDZ basis. The mean signed error (MSE), mean unsigned error (MUE), and standard deviations (SD) are shown at the top of each plot. Reference values are of at least of EOM-CCSDT quality.
S5 Geometries
Geometries for all molecules in the QUEST and CT benchmarks may be found in the Supporting Information from our previous work.[21] The geometries for the large molecule tests may be found below in Angstroms.
Thiopropynal alkane chain
H     17.22333472     -2.23915362     -0.03663289
C     16.36633614     -1.58788420     -0.03067667
C     15.40552368     -0.81011800     -0.02633299
C     14.24962842      0.05208865     -0.01618250
S     14.42365780      1.68894198     -0.01589282
C     12.91978288     -0.68129515     -0.00697016
H     12.92848499     -1.35261850      0.87545130
H     12.91948877     -1.35965812     -0.88389531
C     11.66900178      0.19160285     -0.00424797
H     11.68786240      0.85658715      0.87801085
H     11.68400899      0.85596281     -0.88708278
C     10.38560742     -0.64441549     -0.00112851
H     10.37917591     -1.30863308      0.88500802
H     10.37493287     -1.30843929     -0.88736459
C      9.11714123      0.21356443      0.00190063
H      9.12851765      0.87741843      0.88793853
H      9.12478282      0.87802384     -0.88373285
C      7.82807710     -0.61299393      0.00426540
H      7.82108794     -1.27678016      0.89068617
H      7.81854533     -1.27784227     -0.88133820
C      6.55946725      0.24485440      0.00552339
H      6.56936720      0.90981034      0.89097232
H      6.56686993      0.90870733     -0.88078266
C      5.26949163     -0.58047861      0.00785620
H      5.26181733     -1.24374203      0.89466527
H      5.26007678     -1.24603874     -0.87721558
C      4.00110337      0.27769660      0.00799701
H      4.01047827      0.94294096      0.89328147
H      4.00920427      0.94135501     -0.87849006
C      2.71074150     -0.54706737      0.00967009
H      2.70234819     -1.21027482      0.89651853
H      2.70162926     -1.21274847     -0.87531687
C      1.44253653      0.31138820      0.00899141
H      1.45155074      0.97677192      0.89418982
H      1.45125791      0.97493918     -0.87758547
C      0.15194430     -0.51302222      0.01006396
H      0.14294281     -1.17616866      0.89695349
H      0.14314317     -1.17879214     -0.87486074
C     -1.11611231      0.34565446      0.00865385
H     -1.10742491      1.01115966      0.89377061
H     -1.10685346      1.00909373     -0.87800746
C     -2.40685576     -0.47852214      0.00920465
H     -2.41635790     -1.14161178      0.89613182
H     -2.41536065     -1.14436111     -0.87567130
C     -3.67480369      0.38031421      0.00716275
H     -3.66641268      1.04592432      0.89220630
H     -3.66509941      1.04364994     -0.87957398
C     -4.96564168     -0.44371458      0.00726944
H     -4.97554446     -1.10675676      0.89422771
H     -4.97387240     -1.10960787     -0.87756811
C     -6.23352221      0.41521780      0.00469499
H     -6.22539802      1.08091953      0.88967337
H     -6.22346537      1.07846062     -0.88210844
C     -7.52440638     -0.40873460      0.00443568
H     -7.53461588     -1.07173877      0.89141850
H     -7.53239042     -1.07466925     -0.88037250
C     -8.79224801      0.45024505      0.00142652
H     -8.78434603      1.11602839      0.88634664
H     -8.78191546      1.11340746     -0.88543487
C    -10.08314236     -0.37367391      0.00087930
H    -10.09355970     -1.03665338      0.88787788
H    -10.09090111     -1.03964250     -0.88390493
C    -11.35095771      0.48530969     -0.00246838
H    -11.34321376      1.15117049      0.88239762
H    -11.34040538      1.14840853     -0.88937736
C    -12.64178754     -0.33864485     -0.00322562
H    -12.65220455     -1.00162487      0.88377478
H    -12.64923185     -1.00465307     -0.88798482
C    -13.90965475      0.52018363     -0.00681707
H    -13.90191061      1.18612668      0.87800134
H    -13.89884300      1.18324380     -0.89376604
C    -15.20013217     -0.30380880     -0.00770684
H    -15.21071702     -0.96721258      0.87924697
H    -15.20754489     -0.97026532     -0.89240151
C    -16.46892214      0.55354390     -0.01145302
H    -16.45910676      1.21879769      0.87251008
H    -16.45590538      1.21579763     -0.89762463
C    -17.75044497     -0.28328228     -0.01234868
H    -17.79682534     -0.93188381      0.87947822
H    -17.79355516     -0.93497942     -0.90208078
H    -18.65132508      0.35250411     -0.01510695
Solvated acetone
C    -0.00002055      0.00000337      0.09952080
C     0.00000339      1.28065824     -0.69730806
C     0.00001356     -1.28066977     -0.69722662
O     0.00000782     -0.00000034      1.31378426
H    -0.00000092      2.13624078     -0.02743871
H    -0.00000131     -2.13623626     -0.02743654
H     0.87877589      1.31456782     -1.34406044
H    -0.87877280      1.31458555     -1.34406760
H     0.87874031     -1.31456662     -1.34411046
H    -0.87874537     -1.31458273     -1.34410233
O     3.75699034     -2.39641269     -0.67104103
H     3.29242149     -3.02350603     -0.09471475
H     3.35701473     -2.50844957     -1.55546127
O    -1.46613260     -2.16747491      2.42602027
H    -2.13989653     -1.62494532      2.87737626
H    -0.91436065     -1.49616908      1.97932927
O    -2.06307331     -0.04994330     -2.51987743
H    -2.68550831     -0.77157770     -2.75792350
H    -2.21400256      0.70375335     -3.12322156
O    -0.39749759     -4.58945221      1.57960600
H    -0.83865144     -3.82937019      1.99606011
H    -0.82238260     -4.68696654      0.71065809
O    -0.87506662      1.67241888      3.21074846
H    -0.53527183      1.21738144      2.40374454
H    -0.20922266      1.40137853      3.86557873
O     2.78396403     -0.00098899     -0.14091159
H     2.59235067      0.03438425     -1.10403582
H     3.27811538     -0.83996407     -0.05482333
O     2.50955567      2.03144873      1.67581953
H     3.34166806      2.42923564      1.41080855
H     2.39578488      1.26749806      1.07915836
O     2.14865106     -3.86370678      1.08778891
H     2.19209564     -3.11255424      1.71032446
H     1.27775815     -4.26871810      1.29995701
O    -2.54662755      2.73528028      1.21455956
H    -2.24313392      2.67711470      2.12779301
H    -2.47889014      1.83196632      0.84913988
O    -0.10663065      1.29332862     -5.15697583
H     0.53229011      1.90074111     -4.73844880
H    -0.97449004      1.59052173     -4.84742509
O    -2.35311310      2.25249902     -3.78660554
H    -1.74582387      2.95327039     -3.47146674
H    -3.10410099      2.33966628     -3.18214335
O    -3.14523032      0.32155588     -0.08352288
H    -4.00083304      0.70728830     -0.28848599
H    -2.68628391      0.26287884     -0.95950042
O     2.32984344     -2.69097528     -3.04302740
H     1.57041244     -2.18253034     -3.37076484
H     1.94287277     -3.47480428     -2.59897174
O    -3.09505244     -2.41492926      0.21412972
H    -3.17719438     -1.44710700      0.09444861
H    -2.62643153     -2.50223382      1.06552146
O     1.73807430      4.22122825     -1.38805105
H     2.45929751      3.70586631     -0.98339746
H     1.01865970      4.25490694     -0.73909786
O    -1.48641651     -4.29292014     -0.96878531
H    -2.09784333     -3.64910229     -0.55990415
H    -1.37825985     -3.97860092     -1.88787867
O     1.27574640     -4.63985768     -1.51178453
H     1.73789465     -4.50067543     -0.67203326
H     0.33232945     -4.61673361     -1.27946455
O     0.19449960     -0.89241286     -3.75972991
H     0.14555589     -0.27352558     -4.52049571
H    -0.53492918     -0.56626047     -3.17938972
O    -0.94718565     -3.35391858     -3.48674418
H    -0.55117449     -2.47228705     -3.65946608
H    -0.25420709     -3.98816961     -3.68070520
O     2.44247635      0.27044678     -2.82870740
H     3.13548562     -0.26170003     -3.22683992
H     1.59600545     -0.14758969     -3.12626202
O     1.69809111      2.76967698     -3.70299374
H     2.05249070      1.91041745     -3.40144345
H     1.71093661      3.33815996     -2.90783828
O     1.52925386     -1.80282831      2.77575581
H     1.06571218     -1.22970636      2.13413613
H     0.83003164     -2.27091485      3.24663809
O    -3.01403612     -0.03290087      2.80499922
H    -3.15234958      0.12089592      1.85814871
H    -2.33607449      0.61303398      3.08300678
O     1.55022139      0.72662949      3.90763753
H     1.97403659      1.25818048      3.21103400
H     1.71297858     -0.19647479      3.64967223
O     0.63568174      3.86416901      2.45934525
H     0.01150064      3.22830992      2.86005479
H     1.36819745      3.29966007      2.14806820
O    -3.41714064     -2.27523543     -2.66102162
H    -3.47619917     -2.50866811     -1.72327927
H    -2.70954915     -2.81839502     -3.04042042
O     3.75142779      2.39874438     -0.95363915
H     3.56904846      1.51818981     -0.56632444
H     3.69998361      2.24857566     -1.90255218
O    -0.65647136      4.45095076      0.20501945
H    -0.13703428      4.48419830      1.03618362
H    -1.43184228      3.90844229      0.45792820
O    -0.99609885      4.26182040     -2.55548624
H    -0.85762000      4.42376263     -1.60626952
H    -0.12550104      4.36906667     -2.95032900
O    -3.44394726      2.97425463     -1.37841687
H    -2.80691111      3.63599260     -1.67337555
H    -3.41210445      3.02334141     -0.41450195
Donor-bridge-acceptor charge transfer
C     13.88569453      0.42596960     -0.00000000
H     13.80214752      1.51838629     -0.00000000
C     15.13477199     -0.09201976     -0.00000000
F     15.28098945     -1.43855420     -0.00000000
C     16.36931012      0.68230307     -0.00000000
H     16.26044750      1.77013029     -0.00000000
C     17.63484516      0.21619576     -0.00000000
H     18.50918975      0.87353576     -0.00000000
F     17.95987257     -1.08424358     -0.00000000
C     12.63419683     -0.41586840     -0.00000000
H     12.63471357     -1.08569488      0.88128524
H     12.63471357     -1.08569489     -0.88128524
C     11.35731032      0.42964186     -0.00000000
H     11.35973473      1.09272777      0.88653442
H     11.35973473      1.09272777     -0.88653442
C     10.08042779     -0.41347143      0.00000000
H     10.08183874     -1.07770683      0.88589687
H     10.08183874     -1.07770683     -0.88589687
C      8.80131760      0.42827754      0.00000000
H      8.80127553      1.09260870      0.88597105
H      8.80127553      1.09260869     -0.88597105
C      7.52232239     -0.41379318      0.00000000
H      7.52274423     -1.07825958      0.88584186
H      7.52274423     -1.07825958     -0.88584186
C      6.24268284      0.42732199      0.00000000
H      6.24235899      1.09181366      0.88589363
H      6.24235899      1.09181366     -0.88589363
C      4.96384164     -0.41503441      0.00000000
H      4.96441521     -1.07959099      0.88582386
H      4.96441521     -1.07959099     -0.88582386
C      3.68368466      0.42532130      0.00000000
H      3.68301436      1.08985586      0.88586681
H      3.68301436      1.08985586     -0.88586681
C      2.40520350     -0.41759308      0.00000000
H      2.40605070     -1.08216961      0.88582303
H      2.40605070     -1.08216961     -0.88582303
C      1.12464043      0.42215011      0.00000000
H      1.12368841      1.08670291      0.88585444
H      1.12368841      1.08670291     -0.88585444
C     -0.15353019     -0.42123955      0.00000000
H     -0.15245217     -1.08582128      0.88582424
H     -0.15245217     -1.08582128     -0.88582424
C     -1.43440024      0.41803865      0.00000000
H     -1.43557329      1.08259892      0.88584898
H     -1.43557329      1.08259892     -0.88584898
C     -2.71231486     -0.42573995      0.00000000
H     -2.71105111     -1.09032114      0.88582591
H     -2.71105111     -1.09032114     -0.88582591
C     -3.99341772      0.41318321      0.00000000
H     -3.99476235      1.07774623      0.88584666
H     -3.99476235      1.07774623     -0.88584666
C     -5.27112845     -0.43090252      0.00000000
H     -5.26972556     -1.09547952      0.88582774
H     -5.26972556     -1.09547952     -0.88582774
C     -6.55239575      0.40776698      0.00000000
H     -6.55386460      1.07232922      0.88584688
H     -6.55386460      1.07232922     -0.88584688
C     -7.82995153     -0.43654686      0.00000000
H     -7.82845954     -1.10111554      0.88582985
H     -7.82845954     -1.10111554     -0.88582985
C     -9.11131138      0.40197251      0.00000000
H     -9.11285721      1.06652855      0.88585179
H     -9.11285721      1.06652856     -0.88585178
C    -10.38873073     -0.44252235     -0.00000000
H    -10.38716968     -1.10708305      0.88582618
H    -10.38716968     -1.10708305     -0.88582618
C    -11.67016855      0.39584089      0.00000000
H    -11.67181364      1.06037715      0.88586700
H    -11.67181364      1.06037715     -0.88586700
C    -12.94728798     -0.44905351     -0.00000000
H    -12.94555550     -1.11359752      0.88583384
H    -12.94555550     -1.11359751     -0.88583384
C    -14.22878741      0.38900672     -0.00000000
H    -14.23083524      1.05343330      0.88589108
H    -14.23083524      1.05343330     -0.88589108
C    -15.50541254     -0.45669273     -0.00000000
H    -15.50193862     -1.12156464      0.88565040
H    -15.50193862     -1.12156464     -0.88565040
C    -16.78778519      0.37930436     -0.00000000
H    -16.79500245      1.04342061      0.88698173
H    -16.79500245      1.04342062     -0.88698173
C    -18.06717389     -0.46559851     -0.00000000
H    -18.06532569     -1.12984475      0.88383629
H    -18.06532569     -1.12984475     -0.88383629
N    -19.33029431      0.28206183     -0.00000000
H    -19.32122636      0.91307888      0.80713882
H    -19.32122636      0.91307888     -0.80713882